Obtain the particular solution


1. dy/dx+2y=y^3e^4x, y(0)=1. ans. y^2 (1-2x)=e^-4x



2.dy/dx - y/x=y^5/x^3 ,y(1)=-1. ans. y^4(3-2x^2)=x^4

1
Expert's answer
2021-11-09T15:16:11-0500

1.


"z=y^{1-3}=y^{-2}"

"\dfrac{dz}{dx}=-2y^{-3}\dfrac{dy}{dx}"

"-2y^{-3}\dfrac{dy}{dx}-2y^{-3}(2y)=-2y^{-3}y^3e^{4x}"

"\dfrac{dz}{dx}-4z=-2e^{4x}"

Integration factor


"\mu(x)=e^{-4x}"

"e^{-4x}\dfrac{dz}{dx}-4e^{-4x}z=-2e^{-4x}e^{4x}"

"d(e^{-4x}z)=-2dx"

Integrate


"\int d(e^{-4x}z)=-\int2dx"

"e^{-4x}z=-2x+C"

"y^{-2}=-2xe^{4x}+Ce^{4x}"

"y^2=\dfrac{1}{-2xe^{4x}+Ce^{4x}}"

"y(0)=1"


"1^2=\dfrac{1}{-2(0)e^{4(0)}+Ce^{4(0)}}=>C=1"

"y^2(1-2x)e^{4x}=1"

"y^2(1-2x)=e^{-4x}"

2.


"z=y^{1-5}=y^{-4}"

"\dfrac{dz}{dx}=-4y^{-5}\dfrac{dy}{dx}"

"-4y^{-5}\dfrac{dy}{dx}+4y^{-5}(\dfrac{y}{x})=-4y^{-5}\dfrac{y^5}{x^3}"

"\dfrac{dz}{dx}+4\dfrac{z}{x}=-\dfrac{4}{x^3}"

Integration factor


"\mu(x)=e^{\int(4/x)dx}=x^{4}"

"x^{4}\dfrac{dz}{dx}-4x^{4}\dfrac{z}{x}=-x^{4}\dfrac{4}{x^3}"

"d(x^{4}z)=-4xdx"

Integrate


"\int d(x^{4}z)=-\int4xdx"

"x^{4}z=-2x^2+C"

"z=-2x^{-2}+Cx^{-4}"

"y^{-4}=-2x^{-2}+Cx^{-4}"

"y(1)=-1"


"(-1)^{-4}=-2(1)^{-2}+C(1)^{-4}=>C=3"

"y^{-4}=-2x^{-2}+3x^{-4}"

"y^4(3-2x^2)=x^4"